OpenStudy (anonymous):

(-27x^3(-x^7))/(16x^4) SIMPLIFY

OpenStudy (pfenn1):

\[\frac{-27x^3(-x^7)}{16x^4} \]Remember three things.\[\frac{x^a}{x^b}=x^ax^{-b}\]\[x^ax^b=x^{a+b}\]\[(-x)(-y)=xy\]